A pair of electric charges has a positive potential energy. This implies that:

a) Both charges are positive

b) Both charges are negative

c) The charges have the same sign

d) The charges have opposite signs

I've already determined that D is wrong, so can you please explain to me which selection is right and why? Thank you.
Answer #1

Solution) The electric potential energy is given by

U = Kq1q2/r

Here K = 9×10^(9) Nm^2/C^2

q1 and q2 are charges

r is separation between them

So in calculating electric potential energy we consider sign of the charge also which determines the electric potential energy to be positive or negative .

If both charges are positive then electric potential energy is positive

U = k(+q1)(+q2)/r = positive value

If both charges are negative then also electric potential energy is positive

U = k(-q1)(-q2)/r = positive value

So from above we can say if charges have same sign then electric potential energy is positive

So options a,b,c implies

If charges are of opposite sign then

U = k(+q1)(-q2)/r = negative value